OpenStudy (anonymous):

PLEASE HELP!!!!!!! Find the common solution to the following pair of lines by graphing them on the same axes. x + y = 4 x - y = 0 The lines intersect at (__, __).

OpenStudy (anonymous):

by graphing, or by thinking?

OpenStudy (anonymous):

you have two numbers that are equal since \(x-y=0\) and they add up to \(4\)

OpenStudy (anonymous):

what number plus itself is equal to four?

OpenStudy (anonymous):

2 @satellite73

OpenStudy (anonymous):

yes, so the answer is \((2,2)\)

OpenStudy (anonymous):

oy okey thank you @satellite73

OpenStudy (anonymous):

cause \(2+2=4\) and \(2-2=0\)

OpenStudy (anonymous):

yw